OpenStudy (anonymous):

Evaluate C(n, n). 0 1 n

OpenStudy (anonymous):

this is confusing can you explain ?

OpenStudy (anonymous):

\[ {n\choose n}= \frac{n!}{n!(n-n)!} \]

OpenStudy (anonymous):

In general \[ {n\choose k} = \frac{n!}{k!(n-k)!} \]

OpenStudy (anonymous):

wouldnt be any of those than

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

wouldnt the awnser be any of the multiple chioce awnsers

OpenStudy (anonymous):

No \[ \frac{n!}{n!(n-n)!}= \frac{1}{0!}= 1 \]

OpenStudy (anonymous):

ok